OpenStudy (anonymous):

A bicyclist is riding with speed v(t) 2t + 5 feet/second. What is the distance traveled from t=1 to t=5

OpenStudy (anonymous):

Plug in t=1 and t=1 then do t=5-t=1 That's your answer.

sam (.sam.):

\[v=2t+5\] Velocity is the rate of change of displacement \[\frac{dx}{dt}=2t+5\] \[dx=2t+5dt\] \[\int\limits_{}^{} dx=\int\limits_{}^{}2t+5dt\] \[\large x=[t^2+5t]_1^5 \] \[\large x=50-6=44ft\]